Your financial planner offers you two different investment plans. Plan X is a $14,000 annual perpetuity Plan Y is a 13-year, $20,000 annual annuity. Both plans will make their first payment one year from today At what discount rate would you be indifferent between these two plans? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Discount rate References eBook & Resources Learning Objective: 06-04 How interest rates are quoted (and misquoted). Worksheet Section 6 2 Valuing Level Cash Flows Annuities and Perpetuities Difficulty: Challenge y work
Answer #1

Let r be the discount rate

14000/r=20000/r*(1-1/(1+r)^13)
=>0.7=1-1/(1+r)^13
=>1/0.3=(1+r)^13
=>r=(10/3)^(1/13)-1
=>r=9.7037%